It tells me to solve the quadriatic equation, 9(w-4)^2 +1 = 0
Answer by stanbon(75887)

Solve the quadratic equation, 9(w-4)^2 +1 = 0
-----
Here is one way to do it:
Divide thru by 9 to get:
(w-4)^2 + (1/9) = 0
---
Factor:
[(w-4)-(i/3)]{(w-4)+(i/3)] = 0
---
Solve for "w":
w-4 = i/3 or w-4 = -i/3
---
w = 4+(i/3) or w = 4-(i/3)
=============================
Cheers,
Stan H.